# Asymptotic behaviors of subcritical branching killed spectrally negative Lévy process

### v1

@@ 1.2. Main results
Combining the result above with (1.22), we immediately recover [10, Theorem 1.2] as a corollary.
Organization of the paper: The rest of the paper is organized as follows. In Section 2, we give some results on spectrally negative Lévy processes which will be used in the proofs of our main results. We establish the conditioned limit theorem for Lévy process in Section 3. The proofs of Theorems 1.1 and 1.2 are given in Section 4, and the proof of Theorem 1.4 is given in Section 5.
@@ 2. Preliminaries
In this section, we always assume that $\xi=((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process with Laplace exponent $\Psi$. Recall that $\psi(q):=\sup\{\lambda\geq 0:\Psi(\lambda)=q\}$ is the right inverse of $\Psi$. It is well known, see [11, Section 8], that for any $q\geq 0$,
(2.1) $\mathbf{E}_{0}\left(e^{-q\tau_{x}^{+}}1_{\{\tau_{x}^{+}<\infty\}}\right)=e^{-\psi(q)x}.$
Moreover, since $\xi$ is spectrally negative, we have
@@
The following result on exit probabilities is contained in [11, Theorem 8.1].
@@
(1) For any $x>0$, if $\Psi^{\prime}(0+)>0$,
(2.3) $\mathbf{P}_{x}(\tau_{0}^{-}<\infty)=1-\Psi^{\prime}(0+)W^{(0)}(x),$
and if $\Psi^{\prime}(0+)\leq 0$, $\mathbf{P}_{x}(\tau_{0}^{-}<\infty)=1$.
(2) For any $0<x\leq y$ and $q\geq 0$,
(2.4) $\mathbf{E}_{x}\left(e^{-q\tau_{y}^{+}}1_{\{\tau_{0}^{-}>\tau_{y}^{+}\}}\right)=\frac{W^{(q)}(x)}{W^{(q)}(y)}.$
The following result, which can be found in [11, Lemma 8.4] and [16, Proposition 1], gives the relationship between $W_{c}^{(q)}$ for different values of $q$, $c$, and the asymptotic behavior of $W^{(q)}(x)$ as $x\to\infty$.
@@
For any $x\geq 0$, the function $q\mapsto W^{(q)}(x)$ may be analytically extended to $q\in\mathbb{C}$. Furthermore, for any $q\in\mathbb{C}$ and $c\in\mathbb{R}$ with $\Psi(c)<\infty$, we have
(2.5) $W^{(q)}(x)=e^{cx}W_{c}^{(q-\Psi(c))}(x),\quad x\geq 0,$
where $W_{c}^{(q-\Psi(c))}$ is the scale function of $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$. Furthermore,
The following lemma gives the joint asymptotic behavior of the tail of $\tau_{0}^{-}$ and the Lévy process $\xi$ when $\mathbf{E}_{0}\left[\xi_{1}\right]>0$.
@@
Assume that $\mathbf{E}_{0}(\xi_{1})>0$ and $\sigma^{2}:=\mathbf{E}_{0}(\xi_{1}^{2})<\infty$, then for any $x>0$,
(2.7) $\lim_{t\to\infty}\mathbf{P}_{x}(\tau_{0}^{-}>t)=\mathbf{P}_{x}(\tau_{0}^{-}=\infty)=\Psi^{\prime}(0+)W^{(0)}(x).$
@@
(2.8) $\lim_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)=\mathbf{P}_{x}(\tau_{0}^{-}=\infty)\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$
@@
Note that (2.7) follows immediately from Theorem 2.1 (1). Fix $t>0$, for $m\in(0,t)$, by the Markov property,
(2.9) $\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)$ $\leq\mathbf{P}_{x}\left(\tau_{0}^{-}>m,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)$
(2.10) $=\mathbf{E}_{x}\left[1_{\{\tau_{0}^{-}>m\}}\mathbf{P}_{\xi_{m}}\left(\xi_{t-m}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)\right].$
@@
(2.11) $\mathbf{P}_{z}\left(\xi_{t-m}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)\to\int_{y}^{\infty}\phi_{\sigma^{2}}(z)\mathrm{d}z.$
@@
(2.12) $\limsup_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)\leq\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(z)\mathrm{d}z.$
@@
(2.13) $\mathbf{P}_{x}\left(\tau_{0}^{-}>m,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)\leq\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)+\mathbf{P}_{x}\left(\tau_{0}^{-}\in(m,t]\right).$
@@
(2.14) $\lim_{m\to\infty}\lim_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>m,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)=\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(z)\mathrm{d}z,$
@@
(2.16) $\liminf_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t>\sqrt{t}y\right)\geq\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(z)\mathrm{d}z.$
@@
The following lemma is an important tool for proving Theorem 1.4.
@@
For any $a>0$, $0<x\leq y$ and nonnegative Borel function $h$, we have
(2.17) $\mathbf{E}_{x}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right]=e^{\psi(a)(x-y)}\mathbf{E}_{x}^{\psi(a)}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right].$
Proof.
By [3, Theorem 6], $\{\tau_{y}^{+}<\tau_{0}^{-}\}\cap\{\tau_{y}^{+}<t\}=\{\tau_{y}^{+}\land t<\tau_{0}^{-}\}\cap\{\tau_{y}^{+}\land t<t\}$ is $\mathcal{F}_{\tau_{y}^{+}\land t}$-measurable. For $a>0$, since $e^{-a\tau_{y}^{+}}1_{\{\tau_{y}^{+}=\infty\}}$=0, using (1.23) with $c=\psi(a)$, we have
(2.18) $\mathbf{E}_{x}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right]=\lim_{t\to\infty}\mathbf{E}_{x}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-},\tau_{y}^{+}<t\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right]$
$=\lim_{t\to\infty}\mathbf{E}_{x}^{\psi(a)}\left[e^{-\psi(a)(\xi_{t}-x)+at}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}1_{\{\tau_{y}^{+}<\tau_{0}^{-},\tau_{y}^{+}<t\}}\right]$
$=\lim_{t\to\infty}\mathbf{E}_{x}^{\psi(a)}\left[e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}1_{\{\tau_{y}^{+}<\tau_{0}^{-},\tau_{y}^{+}<t\}}\mathbf{E}_{x}^{\psi(a)}\left[e^{-\psi(a)(\xi_{t}-x)+at}\Big{|}{\mathcal{F}_{\tau_{y}^{+}\land t}}\right]\right].$
(2.19) $\mathbf{E}_{x}^{\psi(a)}\left[e^{-\psi(a)(\xi_{t}-x)+at}\Big{|}{\mathcal{F}_{\tau_{y}^{+}\land t}}\right]=e^{-\psi(a)(\xi_{\tau_{y}^{+}\land t}-x)+a(\tau_{y}^{+}\land t)}=e^{-\psi(a)(y-x)+a\tau_{y}^{+}}.$
(2.20) $\mathbf{E}_{x}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right]$ $=e^{\psi(a)(x-y)}\mathbf{E}_{x}^{\psi(a)}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right].$
This gives the desired result. ∎
3. Conditioned limit theorems for Lévy processes
The purpose of this section is to prove Theorem 3.5, a conditioned limit theorem for spectrally negative Lévy processes. Theorem 3.5 will play an important role in this paper. We make some preparations first. The following result follows from [9, Lemmas 2.12 and 4.1].
@@ 3. Conditioned limit theorems for Lévy processes
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1})=0$ and (H1). Then for any $x>0$ and $a\in(0,\infty]$, it holds that
where $\sigma^{2}:=\mathbf{E}_{0}(\xi_{1}^{2})$. Furthermore, for any $x>0$ and any bounded continuous function $h$ on $(0,\infty)$, it holds that
(3.2) $\lim_{t\to\infty}\sqrt{t}\mathbf{E}_{x}\left[h\left(\frac{\xi_{t}}{\sigma\sqrt{t}}\right)1_{\{\tau_{0}^{-}>t\}}\right]=\frac{2R(x)}{\sqrt{2\pi\sigma^{2}}}\int_{0}^{\infty}\rho(z)h(z)\mathrm{d}z.$
Recall that $\delta$ is the constant in (H1). The following result is [9, Lemma 2.11].
Lemma 3.2.
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1})=0$ and (H1). Then there exists a Brownian motion $W$ with diffusion coefficient $\sigma^{2}$ starting from the origin such that for any $\kappa\in(0,\frac{\delta}{2(2+\delta)})$, there exists a constant $C_{3}(\kappa)>1$ such that for all $t\geq 1$,
@@
The following result is a conditioned limit theorem for Lévy processes with explicit rate of convergence. See [7, Theorem 2.7] for an analogous result for random walks.
Lemma 3.3.
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1})=0$, $\mathbf{E}_{0}(\xi_{1}^{2})=\sigma^{2}$ and (H1). Then one can find a constant $\varepsilon_{0}\in(0,\frac{\delta}{4(2+\delta)})$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{0}(\varepsilon)$ and $C_{4}(\varepsilon)$ such that for any $x,y>0$ and $t>T_{0}(\varepsilon)$,
@@
Let $W$ be the Brownian motion in Lemma 3.2. For any $r>0$ and $\epsilon\in(0,\delta/(2(5+2\delta)))$, define
@@
According to the display below [9, (4.6)], there exist positive constants $c_{1}\in(0,1)$ and $t_{1}(\epsilon)$ such that for $t>t_{1}(\epsilon)$,
@@
(3.12) $I_{1}\leq c_{1}^{K}=e^{-c_{2}[t^{\epsilon}-1]}\leq\frac{c_{2}}{t^{1/2+\epsilon/4}}.$
(ii) Upper bound of $I_{2}$. By part (ii) of the proof of [9, Lemma 4.1], for any $\epsilon\in(0,\delta/(2(5+2\delta)))$, we have
(3.13) $I_{2}\leq\frac{C_{3}(2\epsilon)x}{t^{1+\delta/2-(5+2\delta)\epsilon}}\leq\frac{C_{3}(2\epsilon)x}{t^{\frac{1}{2}+\epsilon/4}},$
@@
(iii) Upper bound of $I_{3}$. Repeating the argument in part (iii) of the proof of [9, Lemma 4.1] leading to [9, (4.8)] and using [6, Lemma 7.7], we can find $\epsilon_{1}>0$ with the property that for any $\epsilon\in(0,\epsilon_{1}\land\delta/(2(5+2\delta)))$ there exists a positive constant $c_{3}(\epsilon)$ such that
@@
(3.15) $\leq\frac{c_{3}(\epsilon)(1+x)}{t^{1+\delta/2-\epsilon(1+\epsilon+\delta/2)}}\leq\frac{c_{3}(\epsilon)(1+x)}{t^{\frac{1}{2}+\epsilon/4}}.$
@@
It follows from [9, (4.13)] that
@@
We claim that there exist positive constants $t_{2}(\epsilon)$ and $c_{4}(\epsilon)$ such that for $t>t_{2}(\epsilon)$,
@@
Thus using the fact that $\rho(z)\leq 1$ for all $z\geq 0$, we get that for any $\epsilon\in(0,\epsilon_{1}\land\delta/(2(5+2\delta)))$ and $t>\max\{t_{i}(\epsilon):3\leq i\leq 5\}$,
@@
For $y>t^{\epsilon^{p}}$, using [6, (7.31)] we get that there exist positive constants $t_{6}(\epsilon)$ and $c_{10}(\epsilon)$ such that for $t>t_{6}(\epsilon)$,
@@
×$\quad\times\sum_{k=1}^{[t^{1-\epsilon}]}\mathbf{E}_{x}\left(\xi_{k}+t^{\frac{1}{2}-2\epsilon};\tau_{0}^{-}>k,\xi_{k}\leq t^{(1-\varepsilon)/2},\tau_{t^{1/2-\epsilon}}^{S,+}=k\right)$

### v2

@@ 1.2. Main results
Remark 4.
The reason we consider spectrally negative Lévy processes here, rather than general Lévy processes, is that the proof of Theorem 1.4 is closely related to the two-sided exit problem. For general Lévy processes, there are no tractable expressions for quantities of interest related to the two-sided exit problem. Combining the result above with (1.37), we immediately recover [12, Theorem 1.2] as a corollary. Profeta [17, Theorem 1.1] proved the following asymptotic behavior of the all-time maximum $\widetilde{M}$ for spectrally negative branching Lévy processes without killing
(1.39) $\mathbb{P}(\widetilde{M}\geq x)\sim\kappa e^{-\psi(\alpha)x},\quad\mbox{ as }x\to\infty,$
under the third-moment condition on the offspring distribution $\{p_{k}\}_{k\geq 0}$, where $\kappa$ is a positive constant. Comparing Theorem 1.4 with (1.39), we observe that the killing barrier does not affect the exponential decay rate of the tail probability of the all-time maximum, it only affects the limits after the same exponential scaling.
1.3. Proof strategies and organization of the paper
The rest of the paper is organized as follows. In Section 2, we give some results on Lévy processes which will be used in the proofs of our main results. We establish the conditioned limit theorem for Lévy processes in Section 3. The proofs of Theorems 1.1 and 1.2 are given in Section 4, and the proof of Theorem 1.4 is given in Section 5.
Now we sketch the main idea of the proof of Theorem $1.$ The main idea for the proof of Theorem 1.2 is similar, and Corollary 1.3 follows from Theorems 1.1 and 1.2. For any $x,t>0$, let
(1.40) $u(x,t):=\mathbb{P}_{x}(\zeta>t).$
In Lemma 4.2, we derive a representation for $u(x,t)$. Lemma 4.3 then establishes a lower bound for $u(x,t)$, while Lemmas 4.4 and 4.5 provide upper bounds for $u(x,t)$ in the cases $\mathbf{E}_{0}(\xi_{1})=0$ and $\mathbf{E}_{0}(\xi_{1})>0$, respectively. Theorem 1.1 (1) and (2) follow immediately from the above lemmas. In the case $\mathbf{E}_{0}(\xi_{1})<0$, a quasi-stationary distribution exists, and the proof technique differs from those used in the previous two cases. The analysis of its asymptotic behavior relies on Theorem 3.5, which establishes a conditioned limit theorem for Lévy processes.
@@ 2. Preliminaries
In this section, we first present some preliminary results for spectrally negative Lévy processes, followed by a result for general Lévy processes. Assume for now that $\xi$ is a spectrally negative Lévy process with Laplace exponent $\Psi$. Then for any $x>0$,
@@
Moreover, it is well known, see [13, Section 8], that for any $x>0$ and $q\geq 0$,
(2.2) $\mathbf{E}_{0}\left(e^{-q\tau_{x}^{+}}1_{\{\tau_{x}^{+}<\infty\}}\right)=e^{-\psi(q)x},$
where $\psi$ is the right inverse of $\Psi$. The following result on exit probabilities is contained in [13, Theorem 8.1].
@@
Assume that $\xi$ is a spectrally negative Lévy process with Laplace exponent $\Psi$. For any $0<x\leq y$ and $q\geq 0$,
(2.3) $\mathbf{E}_{x}\left(e^{-q\tau_{y}^{+}}1_{\{\tau_{0}^{-}>\tau_{y}^{+}\}}\right)=\frac{W^{(q)}(x)}{W^{(q)}(y)},$
where $W^{(q)}$ is the scale function of $\xi$.
The following result, which can be found in [13, Lemma 8.4] and [19, Proposition 1], gives the relationship between $W_{c}^{(q)}$ for different values of $q$, $c$, and the asymptotic behavior of $W^{(q)}(x)$ as $x\to\infty$.
Lemma $2.$
Assume that $\xi$ is a spectrally negative Lévy process with Laplace exponent $\Psi$. For any $x\geq 0$, the function $q\mapsto W^{(q)}(x)$ may be analytically extended to $q\in\mathbb{C}$. Furthermore, for any $q\in\mathbb{C}$ and $c\in\mathbb{R}$ with $\Psi(c)<\infty$, we have
(2.4) $W^{(q)}(x)=e^{cx}W_{c}^{(q-\Psi(c))}(x),\quad x\geq 0,$
where $W_{c}^{(q-\Psi(c))}$ is the scale function of $\xi^{(c)}$. Furthermore,
The following lemma is an important tool for proving Theorem 1.4.
@@
Assume that $\xi$ is a spectrally negative Lévy process with Laplace exponent $\Psi$. For any $a>0$, $0<x\leq y$ and nonnegative Borel function $h$, we have
(2.6) $\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right)=e^{\psi(a)(x-y)}\mathbf{E}_{x}^{\psi(a)}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right).$
Proof.
By Theorem 6 on p16 of [4], $\{\tau_{y}^{+}<\tau_{0}^{-}\}\cap\{\tau_{y}^{+}<t\}=\{\tau_{y}^{+}\land t<\tau_{0}^{-}\}\cap\{\tau_{y}^{+}\land t<t\}$ is $\mathcal{F}_{\tau_{y}^{+}\land t}$-measurable. For $a>0$, since $e^{-a\tau_{y}^{+}}1_{\{\tau_{y}^{+}=\infty\}}$=0, using (1.20) with $c=\psi(a)$, we have
(2.7) $\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right)=\lim_{t\to\infty}\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-},\tau_{y}^{+}<t\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right)$
$=\lim_{t\to\infty}\mathbf{E}_{x}^{\psi(a)}\left(e^{-\psi(a)(\xi_{t}-x)+at}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}1_{\{\tau_{y}^{+}<\tau_{0}^{-},\tau_{y}^{+}<t\}}\right)$
$=\lim_{t\to\infty}\mathbf{E}_{x}^{\psi(a)}\left(e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}1_{\{\tau_{y}^{+}<\tau_{0}^{-},\tau_{y}^{+}<t\}}\mathbf{E}_{x}^{\psi(a)}\left(e^{-\psi(a)(\xi_{t}-x)+at}\Big|{\mathcal{F}_{\tau_{y}^{+}\land t}}\right)\right).$
(2.8) $\mathbf{E}_{x}^{\psi(a)}\left(e^{-\psi(a)(\xi_{t}-x)+at}\Big|{\mathcal{F}_{\tau_{y}^{+}\land t}}\right)=e^{-\psi(a)(\xi_{\tau_{y}^{+}\land t}-x)+a(\tau_{y}^{+}\land t)}=e^{-\psi(a)(y-x)+a\tau_{y}^{+}}.$
(2.9) $\mathbf{E}_{x}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-a\tau_{y}^{+}-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right)$ $=e^{\psi(a)(x-y)}\mathbf{E}_{x}^{\psi(a)}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}h(\xi_{s})\mathrm{d}s}\right).$
This gives the desired result. ∎
The following lemma gives the joint asymptotic behavior of the tail of $\tau_{0}^{-}$ and the Lévy process $\xi$ when $\mathbf{E}_{0}\left(\xi_{1}\right)>0$, and this result holds for general Lévy processes rather than being restricted to the spectrally negative case.
@@
Assume that $\xi$ is a Lévy process such that $\mathbf{E}_{0}(\xi_{1})>0$ and $\sigma^{2}:=\mathbf{E}_{0}(\xi_{1}^{2})<\infty$, then for any $x>0$,
(2.10) $\lim_{t\to\infty}\mathbf{P}_{x}(\tau_{0}^{-}>t)=\mathbf{P}_{x}(\tau_{0}^{-}=\infty)=:q_{x}>0.$
@@
(2.11) $\lim_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)=\mathbf{P}_{x}(\tau_{0}^{-}=\infty)\int_{\frac{y}{\sigma}}^{\infty}\phi(z)\mathrm{d}z.$
@@
Note that (2.10) follows immediately from [3, Proposition 17, p172]. Fix $t>0$, for $m\in(0,t)$, by the Markov property,
(2.12) $\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)$ $\leq\mathbf{P}_{x}\left(\tau_{0}^{-}>m,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)$
(2.13) $=\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>m\}}\mathbf{P}_{\xi_{m}}\left(\xi_{t-m}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)\right).$
@@
(2.14) $\mathbf{P}_{z}\left(\xi_{t-m}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)\to\int_{y}^{\infty}\phi_{\sigma^{2}}(u)\mathrm{d}u.$
@@
(2.15) $\limsup_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)\leq\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(z)\mathrm{d}z.$
@@
(2.16) $\mathbf{P}_{x}\left(\tau_{0}^{-}>m,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)\leq\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)+\mathbf{P}_{x}\left(\tau_{0}^{-}\in(m,t]\right).$
@@
(2.17) $\lim_{m\to\infty}\lim_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>m,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)$
(2.18) $=$ $\lim_{m\to\infty}\lim_{t\to\infty}\mathbf{E}_{x}\left(1_{\{\tau_{0}^{-}>m\}}\mathbf{P}_{\xi_{m}}\left(\xi_{t-m}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)\right)$
(2.19) $=$ $\lim_{m\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>m\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(u)\mathrm{d}u=\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(u)\mathrm{d}u,$
@@
(2.21) $\liminf_{t\to\infty}\mathbf{P}_{x}\left(\tau_{0}^{-}>t,\xi_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t>\sqrt{t}y\right)\geq\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)\int_{y}^{\infty}\phi_{\sigma^{2}}(z)\mathrm{d}z.$
@@
3. Conditioned limit theorems for Lévy processes
The purpose of this section is to prove Theorem 3.5, a conditioned limit theorem for Lévy processes. Theorem 3.5 will play an important role in this paper. We make some preparations first. The following result follows from [11, Lemmas 2.12 and 4.1].
@@ 3. Conditioned limit theorems for Lévy processes
Assume that $\xi$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1})=0$ and (H1). Then for any $x>0$ and $a\in(0,\infty]$, it holds that
where $\sigma^{2}:=\mathbf{E}_{0}(\xi_{1}^{2})$ and $\rho(z)$ denotes the Rayleigh density. Furthermore, for any $x>0$ and any bounded continuous function $h$ on $(0,\infty)$, it holds that
(3.2) $\lim_{t\to\infty}\sqrt{t}\mathbf{E}_{x}\left(h\left(\frac{\xi_{t}}{\sigma\sqrt{t}}\right)1_{\{\tau_{0}^{-}>t\}}\right)=\frac{2R(x)}{\sqrt{2\pi\sigma^{2}}}\int_{0}^{\infty}\rho(z)h(z)\mathrm{d}z.$
Recall that $\delta$ is the constant in (H1) and $\sigma^{2}=\mathbb{E}_{0}(\xi_{1}^{2})$. The following result is [11, Lemma 2.11].
@@
Assume that $\xi$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1})=0$ and (H1). Then there exists a Brownian motion $W$ with diffusion coefficient $\sigma^{2}$ starting from the origin such that for any $\kappa\in(0,\frac{\delta}{2(2+\delta)})$, there exists a constant $C_{3}(\kappa)>1$ such that for all $t\geq 1$,
@@
The following lemma is a conditional limit theorem for Lévy processes. Its proof is similar to that of [11, Lemma 4.1], but it provides a more precise bound. See [8, Theorem 2.7] for an analogous result for random walks.
Lemma $3.$
Assume that $\xi$ is a Lévy process satisfying $\mathbf{E}_{0}(\xi_{1})=0$, $\mathbf{E}_{0}(\xi_{1}^{2})=\sigma^{2}$ and (H1). Then one can find a constant $\varepsilon_{0}\in(0,\frac{\delta}{4(2+\delta)})$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{0}(\varepsilon)$ and $C_{4}(\varepsilon)$ such that for any $x,y>0$ and $t>T_{0}(\varepsilon)$,
@@
Let $W$ be the Brownian motion in Lemma 3.2. For any $r>0$ and $\epsilon\in(0,\delta/(4(5+2\delta)))$, define
@@
According to the display below [11, (4.6)], there exist positive constants $c_{1}\in(0,1)$ and $t_{1}(\epsilon)$ such that for $t>t_{1}(\epsilon)$,
@@
(3.12) $I_{1}\leq c_{1}^{K}=e^{-c_{2}[t^{\epsilon}-1]}\leq\frac{c_{2}}{t^{1/2+\epsilon/8}}.$
(ii) Upper bound of $I_{2}$. By part (ii) of the proof of [11, Lemma 4.1], for any $\epsilon\in(0,\delta/(4(5+2\delta)))$, we have
(3.13) $I_{2}\leq\frac{C_{3}(2\epsilon)x}{t^{1/2+\delta/2-(5+2\delta)\epsilon}}\leq\frac{C_{3}(2\epsilon)x}{t^{\frac{1}{2}+\epsilon/8}},$
@@
(iii) Upper bound of $I_{3}$. Repeating the argument in part (iii) of the proof of [11, Lemma 4.1] leading to [11, (4.8)] and using [7, Lemma 7.7], we can find $\epsilon_{1}>0$ with the property that for any $\epsilon\in(0,\epsilon_{1}\land\delta/(4(5+2\delta)))$ there exists a positive constant $c_{3}(\epsilon)$ such that
@@
(3.15) $\leq\frac{c_{3}(\epsilon)(1+x)}{t^{1+\delta/2-\epsilon(1+\epsilon+\delta/2)}}\leq\frac{c_{3}(\epsilon)(1+x)}{t^{\frac{1}{2}+\epsilon/8}}.$
@@
It follows from [11, (4.13)] that
@@
We claim that there exist positive constants $t_{2}(\epsilon)$, $c_{4}(\epsilon)$ such that for $t>t_{2}(\epsilon)$ and $p\geq 2$ sufficiently large,
@@
Thus using the fact that $\rho(z)\leq 1$ for all $z\geq 0$, we get that for any $\epsilon\in(0,\epsilon_{1}\land\delta/(4(5+2\delta)))$ and $t>\max\{t_{i}(\epsilon):3\leq i\leq 5\}$,
@@
For $y>t^{\epsilon^{p}}$, using [7, (7.31)] we get that there exist positive constants $t_{6}(\epsilon)$ and $c_{10}(\epsilon)$ such that for $t>t_{6}(\epsilon)$,
@@
$\quad\times\sum_{k=1}^{[t^{1-\epsilon}]}\mathbf{E}_{x}\left(\xi_{k}+t^{\frac{1}{2}-2\epsilon};\tau_{0}^{-}>k,\xi_{k}\leq t^{(1-\varepsilon)/2},\tau_{t^{1/2-\epsilon}}^{S,+}=k\right)$